# Closing the Gap Between Text and Speech Understanding in LLMs

### v1

@@ Abstract
Figure 1: SALAD reduces the text-speech understanding gap while requiring over an order of magnitude less training data than competing speech-adapted LLMs. Left: The gap between the performance of the base LLM given text inputs and the performance of the speech-adapted LLM given speech inputs as a function of training dataset size [^2] For Qwen2.5-Omni, we approximate the dataset size in hours based on the reported number of pretraining audio tokens. However, the proportion of this data that is actual speech is not reported.. Right: The performance of the base LLM given text inputs and the performance of the speech-adapted LLM given speech inputs across several language understanding tasks.
@@ 1 Introduction
Large language models (LLMs) have demonstrated impressive capabilities in general knowledge and reasoning, often surpassing specialized systems across a wide range of tasks. This success has motivated efforts to extend LLMs to the speech domain, opening new possibilities for spoken natural interaction. A straightforward approach to extend LLMs to the speech domain is the cascaded pipeline, where automatic speech recognition (ASR) models map speech to text and the LLM is applied to the transcribed text. While effective in preserving text capabilities, cascaded pipelines largely remove speaker and paralinguistic cues essential for natural spoken interaction (Maimon et al., 2025). To address this limitation, recent work has explored end-to-end approaches, adapting text-based LLMs to directly process speech inputs (Tang et al., 2024; Chu et al., 2024; Xie & Wu, 2024; Fang et al., 2024; Nguyen et al., 2024; Défossez et al., 2024). Despite their promise, speech-adapted LLMs struggle with the core requirement of language understanding, consistently under-performing text-based LLMs and even cascaded systems on language understanding tasks (Cuervo & Marxer, 2024; Chen et al., 2024; Cui et al., 2025). We refer to this shortfall as the text–speech understanding gap: the performance drop when a speech-adapted LLM performs a language understanding task in the speech domain compared to the original LLM performing the same task in the text domain. Closing this gap is a crucial step toward building AI systems capable of truly natural spoken interaction.
@@ 2 Preliminaries
Recent end-to-end methods for speech-adapted LLMs typically generate text as an intermediate representation conditioned on speech inputs, and then generate speech conditioned on that text (Xie & Wu, 2024; Défossez et al., 2024; Fang et al., 2024; Xu et al., 2025; KimiTeam et al., 2025). In this work, we focus on generating intermediate text conditioned on speech input as our primary task since it directly reflects language capability, leaving the task of generating speech for future work. Specifically, we aim to build a speech-adapted language model $P_{\theta}$, parameterized by weights $\theta$, that defines a probability distribution over the next text token given a multimodal context. Let $\bm{c}$} denote such a context, which may contain subsequences of text tokens $\bm{w}$} and/or speech representations $\bm{a}$}. For each position $i$, the model predicts the distribution over the next text token $w_{i+1}$ conditioned on ${\bm{c}}_{\leq i}$: $P_{\theta}(w_{i+1}\mid{\bm{c}}_{\leq i})$.
@@ 3.1 Objective
We train our models in a pretraining setup, modeling general sequences without the data templates used in instruction-tuned models. This choice reflects our focus on broad-domain alignment, avoiding restriction to dialogue data and acknowledging the scarcity of speech instruction data. Our data therefore consist of multimodal sequences $\bm{c}$} composed of interleaved speech $\bm{a}$} and text $\bm{w}$} inputs, as in Nguyen et al. (2024). For our training objective, we introduce a variable $\alpha\in[0,1]$ that interpolates between a standard maximum likelihood objective, and a cross-modal distillation objective, similar to that used by Held et al. (2025):
@@ 3.5 Results
Figure 4 reports scaling curves for cross-modal misalignment, forgetting, and average speech performance. Training with NLL (i.e., $\alpha=0$) on narrow-domain speech data (LibriHeavy + Emilia) leads to increasing cross-modal misalignment with scale. Given the strong relation between misalignment and speech performance shown in Figure 3, NLL training on narrow-domain data also yields the weakest results. NLL training leads to greater forgetting of the pretrained text behavior compared to models trained with nonzero $\alpha$ values. This greater forgetting, however, has limited impact on the average text performance (see Appendix A.7).
@@ A.2 Training Details: Analyzing the Text-Speech Gap
The models are trained with the AdamW optimizer (Loshchilov & Hutter, 2019) with a weight decay of $0.1$. We adopt a warmup-stable-decay learning-rate schedule (Hägele et al., 2024), consisting of a linear warmup of $500$ steps followed by a linear decay to zero over the final $20\%$ of training. The peak learning rate is tuned separately for each model size, with distinct values for the language-model backbone and the adapter. We use a learning rate of $10^{-3}$ for the adapter and $5\times 10^{-5}$ for the LLM. All models are trained with a batch size of approximately 1M tokens and a context window of 2048 tokens.
@@ A.3 Training Details: Closing the Text-Speech Gap
For Stage II training of SALAD models, we resume from the last checkpoint before learning-rate decay and continue for 1.9B tokens with a linear decay of the learning rate. An exception is the experiments in Figure 8 (Appendix A.8), where training was resumed from the checkpoint after decay and continued for 950M tokens with a constant learning rate fixed to the post-decay value. These runs were conducted before we identified the setup that yielded stronger results for SALAD models.
A.4 Evaluation Protocol
@@
A.5 Cluster Annotation Pipeline
Phase A: Explain.
Phase B: Judge.
Prompt format.
Listing 1: Explain prompt.
System:
You are an expert at explaining clusters of short texts.
Return STRICT JSON with fields:
User:
"cluster_id": int,
"distance_metric": "cosine" | "l2",
"positives": [ {"id": int, "distance": float, "text": str}, ... ],
"negatives": [ {"id": int, "from_cluster": int,
"distance": float, "text": str}, ... ],
Use distances as prototypicality cues."
Listing 2: Judge prompt.
User:
"categories": [ {"id": int, "title": str, "description": str}, ... ],
"sample": { "text": str },
"instructions": "Pick exactly one best-matching category id."
@@ A.8 Active Selection Analyses
Figure 6: Stage II performance for SALAD-3B with different synthetic data budgets and selectivity $\gamma$ values.
@@
Table 8: Results with mismatched Stage II and evaluation speakers (VoiceBench speaker).
Model MMSU (VoiceBench) OBQA (VoiceBench)
@@
SALAD-3B Stg. I 47.3 14.6 65.5 16.3
SALAD-3B Stg. II 52.4 9.5 74.3 7.5

### v2

@@ Abstract
Figure 1: SALAD reduces the text-speech understanding gap on general reasoning and language understanding benchmarks while requiring over an order of magnitude less training data than competing speech-adapted LLMs. Left: The gap between the performance of the base LLM given text inputs and the performance of the speech-adapted LLM given speech inputs as a function of training dataset size [^2] For Qwen2.5-Omni, we approximate the dataset size in hours based on the reported number of pretraining audio tokens. However, the proportion of this data that is actual speech is not reported.. Right: The performance of the base LLM given text inputs and the performance of the speech-adapted LLM given speech inputs across several language understanding tasks.
@@ 1 Introduction
Large language models (LLMs) have demonstrated impressive capabilities in general knowledge and reasoning, often surpassing specialized systems across a wide range of tasks. This success has motivated efforts to extend LLMs to the speech domain, opening new possibilities for spoken natural interaction. A straightforward approach to extend LLMs to the speech domain is the cascaded pipeline, where automatic speech recognition (ASR) models map speech to text and the LLM is applied to the transcribed text. While effective in preserving text capabilities, cascaded pipelines largely remove speaker and paralinguistic cues essential for natural spoken interaction (Maimon et al., 2025b). To address this limitation, recent work has explored end-to-end approaches, adapting text-based LLMs to directly process speech inputs (Tang et al., 2024; Chu et al., 2024; Xie and Wu, 2024; Fang et al., 2024; Nguyen et al., 2024; Défossez et al., 2024). Despite their promise, speech-adapted LLMs struggle with the core requirement of language understanding, consistently under-performing text-based LLMs and even cascaded systems on language understanding tasks (Cuervo and Marxer, 2024; Chen et al., 2024; Cui et al., 2025). We refer to this shortfall as the text–speech understanding gap: the performance drop when a speech-adapted LLM performs a language understanding task in the speech domain compared to the original LLM performing the same task in the text domain. Closing this gap is a crucial step toward building AI systems capable of truly natural spoken interaction.
@@ 2 Preliminaries
Recent end-to-end methods for speech-adapted LLMs typically generate text as an intermediate representation conditioned on speech inputs, and then generate speech conditioned on that text (Xie and Wu, 2024; Défossez et al., 2024; Fang et al., 2024; Xu et al., 2025; KimiTeam et al., 2025). In this work, we focus on generating intermediate text conditioned on speech input as our primary task since it directly reflects language capability, leaving the task of generating speech for future work. Specifically, we aim to build a speech-adapted language model $P_{\theta}$, parameterized by weights $\theta$, that defines a probability distribution over the next text token given a multimodal context. Let $\bm{c}$} denote such a context, which may contain subsequences of text tokens $\bm{w}$} and/or speech representations $\bm{a}$}. For each position $i$, the model predicts the distribution over the next text token $w_{i+1}$ conditioned on ${\bm{c}}_{\leq i}$: $P_{\theta}(w_{i+1}\mid{\bm{c}}_{\leq i})$.
@@ 3.1 Objective
We train our models in a pretraining setup, modeling general sequences without the data templates used in instruction-tuned models. This choice reflects our focus on broad-domain alignment, avoiding restriction to dialogue data and acknowledging the scarcity of speech instruction data. Our data therefore consist of multimodal sequences $\bm{c}$} composed of interleaved speech $\bm{a}$} and text $\bm{w}$} inputs, as in Nguyen et al. (2024). For our training objective, we introduce a variable $\alpha\in[0,1]$ that interpolates between a standard maximum likelihood objective, and a cross-modal distillation objective, similar to that used by Wang et al. (2024); Held et al. (2025):
@@ 3.5 Results
Figure 4 reports scaling curves for cross-modal misalignment, forgetting, and average speech performance. Training with NLL (i.e., $\alpha=0$ in Equation 4) on narrow-domain speech data (LibriHeavy + Emilia) leads to increasing cross-modal misalignment with scale. Given the strong relation between misalignment and speech performance shown in Figure 3, NLL training on narrow-domain data also yields the weakest results. NLL training leads to greater forgetting of the pretrained text behavior compared to models trained with nonzero $\alpha$ values. This greater forgetting, however, has limited impact on the average text performance (see Appendix A.7).
@@ 4.3 Results
5 Related Work
Cross-modal transfer in speech-adapted LLMs.
A large body of work has focused on transferring the general language capabilities of text LLMs to the speech domain. Deng et al. (2025); Wang et al. (2023); Tang et al. (2024); Chu et al. (2023; 2024); Hu et al. (2024) adapted text LLMs to process speech through multi-task supervised learning, combining tasks like speech recognition with speech instruction-following data to promote general-domain transfer. Other works, such as Fathullah et al. (2024); Wang et al. (2024); Held et al. (2025), addressed the scarcity of broad-domain supervised data by deriving rich supervision from unsupervised paired text–speech datasets via cross-modal distillation. Related to our method with $\alpha=1.0$ in Equation 4, Wang et al. (2024); Held et al. (2025) showed that dense supervision—matching the full output distribution—in cross-modal distillation can outperform prior supervised-learning-based approaches. Several works (Wang et al., 2021; Held et al., 2025; Tseng et al., 2025) further promoted transfer through specialized representation alignment modules, typically transformer encoders with downsampling and query-based cross-attention, and Wang et al. (2024); Tseng et al. (2025) additionally matched sampling rates across modalities. While these works demonstrated varying degrees of cross-modal knowledge transfer, none matched the performance of text models of equivalent capacity on reasoning- and knowledge-intensive LLM benchmarks. Among open-training-recipe models, SALAD reduces the text–speech understanding gap by 11.7% over the next-best model and rivals the strongest closed-recipe system—within 1.2%—while using over an order of magnitude less training data and maintaining expressive, streaming-friendly speech representations. Our analysis of forgetting and misalignment also provides actionable insights into the factors driving the text–speech gap and the training dynamics of such models.
Domain matching text–speech pretraining.
Cuervo and Marxer (2024) demonstrated the benefits for speech LM training of using a TTS-synthesized spoken version of a small pretraining corpus designed to promote essential language capabilities. Zeng et al. (2025) synthesized an LLM pretraining corpus (FineWeb-Edu, corresponding to ${\mathbb{D}}_{\text{web}}$ in our study) for interleaved text–speech language modeling. Our active selection approach enables the use of far less synthetic speech by employing it adaptively to cover domain gaps in natural speech corpora. Moreover, we show that domain matching should be paired with cross-modal distillation to achieve general-domain alignment.
@@ A.2 Training Details: Analyzing the Text-Speech Gap
The models are trained with the AdamW optimizer (Loshchilov and Hutter, 2019) with a weight decay of $0.1$. We adopt a warmup-stable-decay learning-rate schedule (Hägele et al., 2024), consisting of a linear warmup of $500$ steps followed by a linear decay to zero over the final $20\%$ of training. The peak learning rate is tuned separately for each model size, with distinct values for the language-model backbone and the adapter. We use a learning rate of $10^{-3}$ for the adapter and $5\times 10^{-5}$ for the LLM. All models are trained with a batch size of approximately 1M tokens and a context window of 2048 tokens.
@@ A.3 Training Details: Closing the Text-Speech Gap
For Stage II training of SALAD models, we resume from the last checkpoint before learning-rate decay and continue for 1.9B tokens with a linear decay of the learning rate. An exception is the experiments in Figure 8 and Table 8 (Appendix A.8), where training was resumed from the checkpoint after decay and continued for 950M tokens with a constant learning rate fixed to the post-decay value. These runs were conducted before we identified the setup that yielded stronger results for SALAD models.
During initial Stage II experiments, we observed that clusters with high misalignment often corresponded to domains such as academic citations (including DOIs and URLs) or non-English content, which produced hard-to-parse TTS artifacts. These clusters were inherently difficult to align not necessarily because of domain gaps, but because of mismatches between the TTS-generated speech and the ground-truth text. Consequently, we filtered out the top 15% of clusters (19 clusters for our 128-cluster setup) with the highest character error rate (CER), measured by transcribing the generated speech with Whisper-v3-large and comparing it to the ground-truth text. After removing these TTS-adverse clusters, we observed no meaningful correlation between CER and misalignment, and therefore attribute the remaining misalignment to domain gaps.
A.4 Cluster Annotation Pipeline
Phase A: Explain.
Phase B: Judge.
Prompt format.
Listing 1: Explain prompt.
System:
You are an expert at explaining clusters of short texts.
Return STRICT JSON with fields:
User:
"cluster_id": int,
"distance_metric": "cosine" | "l2",
"positives": [ {"id": int, "distance": float, "text": str}, ... ],
"negatives": [ {"id": int, "from_cluster": int,
"distance": float, "text": str}, ... ],
Use distances as prototypicality cues."
Listing 2: Judge prompt.
User:
"categories": [ {"id": int, "title": str, "description": str}, ... ],
"sample": { "text": str },
"instructions": "Pick exactly one best-matching category id."
A.5 Evaluation Protocol
@@ A.8 Active Selection Analyses
Figure 6: Stage II performance for SALAD-3B with different synthetic data budgets and selectivity $\gamma$ values ($K=128$ clusters).
@@
Table 8: Performance (Accuracy, $\%$) of the SALAD-3B model after Stage II training with active selection (synthetic data budget of 5% of the natural speech data budget, $\gamma=5.0$) across different number of clusters $K\in\{64,128,256\}$.
$K$ SC MMSU OBQA HellaSwag ARC-C PIQA Avg
64 74.8 52.4 74.3 68.9 78.4 78.1 71.1
128 74.5 52.5 75.8 68.7 78.4 78.2 71.4
256 74.1 51.5 72.1 68.6 77.8 78.5 70.4
@@
Is active selection sensitive to the number of clusters used?
We present in Table 8 the performance for different values of the number of clusters $K\in\{64,128,256\}$, using $\gamma=5$ and a Stage II synthetic data budget of 5% of the natural speech dataset size (LibriHeavy + Emilia). We find that varying $K$ has only a negligible effect on the overall average performance.
@@
Table 9: Impact of mismatched Stage II and evaluation speakers for SALAD-3B on VoiceBench MMSU and OpenBookQA.
Evaluation speaker MMSU OBQA
@@
Kokoro TTS; af-heart voice (Original) 52.5 9.4 76.7 5.1
Kokoro TTS; af-bella voice 53.0 8.9 76.0 5.8
VoiceBench Google TTS; male voice 52.4 9.5 74.3 7.5
A.9 Open-ended generation evaluations
SALAD is a pre-training method, and the standard way to evaluate pre-trained LLMs is through multiple-choice benchmarks, where likelihood-based accuracies can be computed—such as the benchmarks used throughout the main body of this paper. These evaluations directly align with the pre-training objective and are therefore well suited for base models. In contrast, the open-ended generation benchmarks commonly used for instruction-tuned LLMs rely on controlled, instruction-following behavior, and are not applicable to our base model, which does not reliably follow formatting or prompting constraints.
Nevertheless, to obtain a coarse sense of the generation quality of SALAD-trained models, we follow prior work on evaluating pre-trained LMs (Lakhotia et al., 2021; Maimon et al., 2025a; Eldan and Li, 2023) and compute three generation-focused metrics on samples generated from 300 LibriHeavy test-clean prefixes. For all generations, we use top-$k$ sampling with $k=250$ and temperature 0.7.
  • GenPPL (Lakhotia et al., 2021; Maimon et al., 2025a): perplexity of generated text under an independent reference LM (SmolLM-1.7B), assessing fluency;
  • AutoBLEU (Lakhotia et al., 2021): a self-repetition metric quantifying diversity and penalizing over-repetitive behavior typical of base models; and
  • LLM-as-a-judge scores (Eldan and Li, 2023), evaluating coherence, relevance, creativity, and fluency on a 1–10 scale. The prompt is shown in Listing 3. We use Claude 3.7 Sonnet with greedy sampling as judge.
To determine whether SALAD training affects generation quality relative to the text-only backbone, we report these metrics for SALAD-3B and its Qwen2.5–3B backbone model. For SALAD–3B, we evaluate generations conditioned on (i) text prompts, (ii) natural speech prompts, and (iii) synthetic speech prompts produced using the same setup as in our main evaluations—Kokoro TTS with the af-heart voice. All results, together with Welch’s unpaired two-sample p-values computed against the Qwen2.5–3B backbone, are shown in Table 10. Across prompting modalities no metric shows any statistically significant difference relative to the Qwen2.5–3B backbone (Holm-adjusted Mann–Whitney tests), and we conclude no clear degradation in generation quality.
Table 10: Generation-quality evaluation for Qwen2.5–3B and SALAD–3B. Metrics include GenPPL, AutoBLEU, and LLM-as-a-judge scores (1–10 scale). For GenPPL and AutoBLEU we report the median (Q1–Q3), and Holm–Bonferroni adjusted Mann–Whitney $p$-values against the Qwen2.5–3B baseline. Speech prompts for SALAD–3B use natural LibriHeavy speech or synthetic speech generated with Kokoro TTS af-heart.
GenPPL AutoBLEU LLM-as-a-judge
Median (Q1–Q3) $p$-value Median (Q1–Q3) $p$-value Median (Q1–Q3) $p$-value
Qwen2.5–3B (baseline) $3.519\ (1.608\text{--}11.540)$ $0.257\ (0.195\text{--}0.302)$ $7.50\ (6.75\text{--}8.25)$
SALAD-3B
 Given text $3.105\ (1.504\text{--}7.660)$ 0.651 $0.269\ (0.205\text{--}0.339)$ 0.126 $7.50\ (7.00\text{--}8.25)$ 1.000
 Given natural speech $4.103\ (1.793\text{--}9.233)$ 1.000 $0.247\ (0.190\text{--}0.323)$ 1.000 $7.50\ (7.00\text{--}8.25)$ 1.000
 Given synthetic speech $3.855\ (1.960\text{--}9.028)$ 1.000 $0.245\ (0.205\text{--}0.321)$ 1.000 $7.50\ (7.00\text{--}8.25)$ 1.000
Listing 3: LLM-as-a-judge prompt.
System:
You are an expert literary critic and linguist, specialising in the evaluation of AI-generated creative text.
You MUST return ONLY a valid JSON object with this exact schema:
"coherence": <int 1-10>,
"relevance": <int 1-10>,
"creativity": <int 1-10>,
"fluency": <int 1-10>,
"coherence_just": "<brief explanation>",
"relevance_just": "<brief explanation>",
"creativity_just": "<brief explanation>",
"fluency_just": "<brief explanation>",
"summary": "<one paragraph synthesis>"
Do NOT wrap the JSON in markdown backticks. Do NOT add any extra keys.
User:
INSTRUCTIONS FOR EVALUATION SCOPE:
The completions come from a base LLM truncated at a fixed length given a small book snippet as prefix, therefore:
1. Identify the core completion portion of the generated completion. E.g. ignore any trailing incomplete sentences or clearly off-topic text product of the LLM effectively switching to a different document.
2. Your analysis must stop at the end of the last complete sentence of this core completion.
3. Ignore any trailing incomplete sentences, loops, meta-commentary or tokens like <|endoftext|>.
SPECIAL CONSIDERATIONS FOR THIS EVALUATION (included ONLY for audio_bin / tts_mimi modalities):
* Punctuation Leniency: The prompt arrived via audio; if the generation continues the last sentence of the prompt, do not penalise sentence fusion.
* Name Leniency: Accept minor variations in proper nouns if the characters role and gender stays consistent.
Story Prompt:
> {STEM_PROMPT}
Generated Completion:
> {GENERATION}
Please provide the JSON rubric now.